OpenStudy (anonymous):

Given the function rule f(x)=x^2-3x+2 what is the output for f(-2) ??? medal

OpenStudy (anonymous):

@iGreen

OpenStudy (igreen):

Just plug in -2 for 'x'

OpenStudy (igreen):

\(\sf f(x) = x^2 - 3x + 2\) \(\sf f(-2) = -2^2 - 3(-2) + 2\)

OpenStudy (anonymous):

So.. 12?

OpenStudy (igreen):

Yep, you got it.

OpenStudy (anonymous):

That was so easy thank you! :D @iGreen

OpenStudy (igreen):

Np